# Profiling stellar environments of gravitational wave sources

### v1

@@ Abstract
Results.— We present results for three CBC environments: vicinity of SMBH (Keplerian potential), GC (Plummer Plummer (1911) profile), NSC (Bahcall-Wolf Hoang et al. (2018) profile); three orbital shapes: circular, elliptic and effective Keplerian [^5] What we can call an “effective Keplerian orbit” is a bound orbit that is not necessarily closed.; and four single-detector configurations: LIGO at A+ sensitivity, Einstein Telescope (ET), DECIGO, and LISA. While circular orbits are possible in all attractive potentials, closed elliptical orbits are only possible in attractive potentials with $1/r$ and $1/r^{2}$ profiles. Moreover, only those orbits with a changing $r$ will be sensitive to the steepness of the potential profile. Thus, for Bahcall-Wolf and Plummer potentials, we only consider effective Keplerian orbits by evolving both $r$ and $\vartheta$ over the duration of the GW signal within the frequency band of the detectors.
@@
For binaries in a circular orbit around an SMBH with Schwarzschild radius $R_{\rm s}$, the kinematic parameters along our LOS depend on the mass of the SMBH ($M_{\rm SMBH}$), the radius of the circular orbit ($R$), and $\cos\vartheta$ (see Figure 1). We use the first three derivatives of LOSV, viz. LOSA, LOSJ, and LOS snap (LOSS), to uniquely determine these quantities [^6] Note that $M_{\rm SMBH}$ and $R$ will still be degenerate with $\iota_{\mathrm{out}}$ and this degeneracy can not be broken even if we consider one more derivative. Similar degeneracies exist between the environment/orbit parameters, and $\iota_{\mathrm{out}}$, for all scenarios considered in this work. We set $\sin\iota_{\mathrm{out}}=1$ throughout; thus, the relative errors in the inference of environment and orbit parameters should be thought of as lower limits..
@@
Figure 2 shows the relative errors in the measurements of $M_{\rm SMBH}$ (top panel) and $R$ (bottom panel) over a $M_{\rm SMBH}-R$ grid. We denote the inference of a parameter as “successful” if the relative r.m.s. error is less than unity. We find that:
@@
    At A+ sensitivity, we can successfully extract the mass of a $M_{\mathrm{SMBH}}=10^{5}\,M_{\odot}$ SMBH and localise a $5-1.4\,M_{\odot}$ binary having optimal SNR [^7] Note that for all of the cases considered in this work, the SNRs quoted are for face-on inner orbits. $\sim 103$ corresponding to a luminosity distance $D_{L}=100\,{\rm Mpc}$, out to $R\sim 200\,R_{s}$.
@@
    With ET, we can successfully extract the mass of the same SMBH and localise a $30-30\,M_{\odot}$ binary having optimal SNR $\sim 6610$ corresponding to $D_{L}=100\,{\rm Mpc}$, out to $R\sim 250\,R_{s}$, and within a precision of $\sim 10\%$ up to $R\sim 100\,R_{s}$.
@@
    We can successfully infer the mass of an SMBH as large as $10^{8}\,M_{\odot}$, and localise a $100-100\,M_{\odot}$ binary having optimal SNR $\sim 7624\;(10)$ corresponding to $D_{L}=1{\rm Gpc}$, out to $\sim 10^{5}\,R_{s}$ ($\sim 2\times 10^{4}\,R_{s}$) with DECIGO (LISA).
Furthermore, for tight orbits around low mass black holes we can extract these quantities with sub-percent precision. All results in this analysis assume $\cos\vartheta$ to be 0.9. In the Supplemental Material Tiwari et al. (2024a), we show that the precision of the inferred environment and orbit parameters do not depend strongly on the value of $\cos\vartheta$.
For binaries in an elliptical orbit around an SMBH, the kinematic parameters along our LOS depend on $M_{\rm SMBH}$, semi-major axis ($R$), eccentricity ($e$), and $x\equiv R/r$ where $r$ is distance of the CoM of the binary from the gravitating center, at $t_{c}$. Therefore, together with LOSA, LOSJ, and LOSS, we will need one more derivative of LOSV: LOS crackle (LOSC), to uniquely determine these quantities (see Tiwari et al. (2024a) for the expressions of kinematic parameters).
@@
    With LIGO at A+ sensitivity, we can successfully extract the mass of a $10^{5}\,M_{\odot}$ SMBH and localise a $5-1.4\,M_{\odot}$ binary having optimal SNR $\sim 103$ corresponding to $D_{L}=100\,{\rm Mpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 250\,R_{s}$.
@@
    With ET, we can successfully extract the mass of the same SMBH and localise a $30-30\,M_{\odot}$ binary having optimal SNR $\sim 6610$ corresponding to $D_{L}=100\,{\rm Mpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 500\,R_{s}$, and within a precision of $\sim 20\%$ up to $R\sim 200\,R_{s}$.
@@
    We can successfully infer the masses of SMBHs as massive as $10^{8}\,M_{\odot}$, and localise a $100-100\,M_{\odot}$ binary having optimal SNRs $\sim 7624$ and $\sim 10$, corresponding to $D_{L}=1{\rm Gpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 5\times 10^{4}\,R_{s}$ with DECIGO and LISA.
Note that the chosen value of $x$ is a random draw from a uniform distribution between $\frac{1}{1-e},\frac{1}{1+e}$. We also find that, for certain portions of the parameter space, we can achieve sub-percent precision in DECIGO (see lower left regions of the parameter space in Figure 3).
For binaries in NSCs with a Bahcall-Wolf (BW) potential profile, having a massive black hole (MBH) at their center, the kinematic parameters projected along the LOS will depend on $M_{\rm MBH}$, $r$, steepness of the potential profile ($\alpha$), radial ($v_{r}$) and angular ($\Omega$) speeds of the CBC’s CoM at coalescence, and $\cos\vartheta$. Hence, together with LOSA, LOSJ, LOSS, and LOSC, we will need two more derivatives of LOSV to break all degeneracies among the inferred parameters.
@@
    With DECIGO, we can extract the mass of a $10^{6}\,M_{\odot}$ MBH within a precision of $\sim 10\%$ together with $\alpha$, and localise a $30-30\,M_{\odot}$ binary having optimal SNR $\sim 2794$ corresponding to $D_{L}=1\,{\rm Gpc}$, with a sub-percent precision out to $\sim 3\times 10^{5}\,R_{s}$.
@@
    With LISA, we can extract the mass of a $10^{6}\,M_{\odot}$ MBH within a precision of $\sim 25\%$ together with $\alpha$ and localise a $100-100\,M_{\odot}$ binary having optimal SNR $\sim 10$ corresponding to $D_{L}=1\,{\rm Gpc}$, with a sub-percent precision out to $\sim 4\times 10^{4}\,R_{s}$.
@@
It is clear from Figure 10 that we can extract $M_{\rm MBH}$ and $r$ with sub-percent precision for certain parts of this parameter space, for both DECIGO and LISA. However, we cannot do so for $\alpha$ with LISA, for the fiducial CBC we consider. This can be understood as follows. Larger magnitudes of kinematic parameters increase our ability to constrain environment parameters. Getting closer to the MBH does indeed increase these magnitudes; however, the MBH potential starts to dominate over the potential with exponent $\alpha$. As a result, the precision of $\alpha$ gets suppressed. There is therefore a sweet spot in the $M_{\rm MBH}-r$ parameter space where the precision on the inference of $\alpha$ is optimized. This spot changes, depending on intrinsic parameters of the CBC, including its orbital parameters [^8] Figure 10 does not clearly show this sweet spot because it lies within the highly-relativistic domain where our analysis is not carried out..
@@
Figure 4: Effective Keplerian outer-orbit in an NSC with Bahcall-Wolf profile: The precision of the extraction of parameters over a grid of MBH masses $M_{\rm MBH}$ and merger distance $r$ for a binary system of component masses (from left to right): $30-30M_{\odot}$ and $100-100M_{\odot}$ at $1{\rm Gpc}$ in DECIGO and LISA, respectively, for $\alpha=1.75$ and $\cos\vartheta=0.9$ with radial and tangential velocities: $v_{r}=250{\rm km/s}$ and $r\Omega=100{\rm km/s}$ [^10] These numbers are consistent with observed proper motions of stars in the Galactic NSC (see Figure 6 of Schödel et al. (2009))., respectively. Top Panel: Relative error in the mass of the MBH. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm MBH}}{M_{\rm MBH}}$. Middle Panel: Relative error in the distance of the binary from the NSC’s center. The ‘dashed’ lines are the lines of constant $\frac{\Delta r}{r}$. Bottom Panel: Relative error in the steepness parameter $\alpha$. The ‘dotted’ lines are the lines of constant $\frac{\Delta\alpha}{\alpha}$. The pink regions, the blue lines, and the arrows have the same meaning as in Figure 2.
@@
For binaries in GCs with a Plummer potential Plummer (1911); Aarseth et al. (1974) profile, kinematic parameters along our LOS will depend on the total mass of the cluster ($M_{0}$), $r$, Plummer radius ($a_{\mathrm{p}}$), $v_{r}$, $\Omega$, and $\cos\vartheta$. Therefore, similar to the BW case, we will need six derivatives of LOSV to uniquely determine parameters. However, the Plummer profile being shallower than the BW profile, recovering LOSC and higher time derivatives of LOSV is not possible for the range of parameters we consider. Hence, we only evaluate the LOSA, LOSJ, and LOSS (see Tiwari et al. (2024a) for the expressions) when constructing the GW phase modulations.
@@
Figure 5 shows the relative errors in the measurements of $M_{0}$ (top panel), $r$ (middle panel), and $a_{\mathrm{p}}$ (bottom panel) over a grid of $M_{0}$ and $r$ for $a_{\rm p}=0.1\,{\rm pc}$. We find that with DECIGO, we can successfully extract $M_{0}$, $a_{\rm p}$, and localise a $30-30\,M_{\odot}$ binary having optimal SNR $\sim 27940$ corresponding to $D_{L}=100\,\,{\rm Mpc}$, for total cluster masses greater than $\sim 3\times 10^{5}\,M_{\odot}$. Even though we can successfully constrain $M_{0}$, $r$, and $a_{\mathrm{p}}$ out to $r\sim 0.4\mathrm{pc}$, it still requires the cluster to be as heavy as $10^{7}\,M_{\odot}$. Note that with LISA, LOSJ is measurable for a very limited portion of the parameter space we consider. Therefore, we do not show the corresponding results.
@@
Figure 5: Effective Keplerian outer-orbit in a GC with Plummer profile: The precision of the extraction of parameters over a grid of cluster masses $M_{0}$ and merger distance $r$ for a binary system of component masses $30-30\,M_{\odot}$ at $100\mathrm{Mpc}$ in DECIGO with $\cos\vartheta=0.25$. Left Panel: Relative error in the total mass of the cluster. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{0}}{M_{0}}$. Middle Panel: Relative error in the distance of the binary from the GC’s center. The ‘dashed’ lines are the lines of constant $\frac{\Delta r}{r}$. Right Panel: Relative error in the Plummer radius where $a_{\rm p}=0.1{\rm pc}$. The ‘dotted’ lines are the lines of constant $\frac{\Delta a_{\mathrm{p}}}{a_{\mathrm{p}}}$. The pink regions and the arrows have the same meaning as in Figure 2.

### v2

@@ III Results
We present results for three CBC environments: vicinity of supermassive black hole (SMBH, Keplerian potential), globular cluster (GC, Plummer Plummer (1911) profile), nuclear star cluster, (NSC, Bahcall-Wolf Hoang et al. (2018) profile); three orbital shapes: circular, elliptic and effective Keplerian — a bound orbit that is not necessarily closed; and four single-detector configurations: LIGO at A+ sensitivity, ET, DECIGO, and LISA. We do not consider outer orbits around stellar mass BHs, which, for numerous configurations, could be chaotic. Such systems are analyzed in Samsing et al. (2024); Hendriks et al. (2024) by focusing on the phase-shifts due to the Rømer delay. While circular orbits are possible in all attractive potentials, closed elliptical orbits are only possible in attractive potentials with $1/r$ and $1/r^{2}$ profiles, where $r$ is the radial coordinate in the orbital plane. Moreover, only those orbits with a changing $r$ will be sensitive to the steepness of the potential profile. Thus, for Bahcall-Wolf (BW) and Plummer potentials, we only consider effective Keplerian orbits by evolving both $r$ and $\vartheta$ (angular coordinate) over the duration of the GW signal within the frequency band of the detectors. Additionally, we do not consider the CBC in GC and NSC environments for LIGO at A+ sensitivity and ET. This is because the signal duration will be very small, and for GC case kinematic parameters will be too small to be measurable; while for NSC case even though kinematic parameters can be measurable if the CBC is very close to the centre, the potential of SMBH will be dominating and hence the scenario will be similar to the CBC around an SMBH case.
Figure 3: Eccentric outer-orbit around an SMBH: The precision of the extraction of parameters over a grid of SMBH masses $M_{\rm SMBH}$ and semi-major axes $R$ of the outer orbit for a binary system of component masses (from left to right) considered in Figure 2 with $e=0.5$ and $\approx 1.25. xx$ was drawn randomly from a uniform distribution between $\frac{1}{1-e}$ and $\frac{1}{1+e}$. Top Panel: Relative error in the extraction of mass of SMBHs. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm SMBH}}{M_{\rm SMBH}}$. Bottom Panel: Relative error in the extraction of the location of binary in the outer orbit. The ‘dotted’ lines are the lines of constant $\frac{\Delta x}{x}$. The pink regions, the blue lines, and the arrows have the same meaning as in Figure 2.
III.1 Circular Orbit around an SMBH
@@
For binaries in a circular orbit around an SMBH with Schwarzschild radius $R_{\rm s}$, the kinematic parameters along our LOS depend on the mass of the SMBH ($M_{\rm SMBH}$), the radius of the circular orbit ($R$), and $\cos\vartheta$ (see Figure 1). We use the first three derivatives of LOSV, viz. LOSA, LOSJ, and LOS snap (LOSS), to uniquely determine these quantities. $M_{\rm SMBH}$ and $R$ will still be degenerate with $\iota_{\mathrm{out}}$ and this degeneracy can not be broken, even if we consider one more derivative. Similar degeneracies exist between environment/orbit parameters and $\iota_{\mathrm{out}}$ for all scenarios considered in this work. We set $\sin\iota_{\mathrm{out}}=1$ throughout; thus, the relative errors in the inference of environment and orbit parameters should be thought of as lower limits. An alternative would have been to assume an uninformative prior on the inclination, which would increase the uncertainty in the inference of environment and orbit parameters by a factor of a few.
@@
Figure 2 shows the relative errors in the measurements of $M_{\rm SMBH}$ (top panel) and $R$ (bottom panel) over a $M_{\rm SMBH}-R$ grid. We denote the inference of a parameter as “successful” if the relative r.m.s. error is less than unity. For completeness, we also give the optimal signal-to-noise ratios (SNRs) (see II.4) of the GW events we consider. We find that:
@@
    At A+ sensitivity, we can successfully extract the mass of a $M_{\mathrm{SMBH}}=10^{5}\,{\rm M}_{\odot}$ SMBH and localise a $5-1.4\,{\rm M}_{\odot}$ binary having optimal SNR $\sim 145$ corresponding to a luminosity distance $D_{\rm L}=100\,{\rm Mpc}$, out to $R\sim 200\,R_{\rm s}$.
@@
    With ET, we can successfully extract the mass of the same SMBH and localise a $30-30\,{\rm M}_{\odot}$ binary having optimal SNR $\sim 8248$ corresponding to $D_{\rm L}=100\,{\rm Mpc}$, out to $R\sim 250\,R_{\rm s}$, and within a precision of $\sim 10\%$ up to $R\sim 100\,R_{\rm s}$.
@@
    We can successfully infer the mass of an SMBH as large as $10^{8}\,{\rm M}_{\odot}$, and localise a $100-100\,{\rm M}_{\odot}$ binary having optimal SNR $\sim 10782\;(14)$ corresponding to $D_{\rm L}=1\,{\rm Gpc}$, out to $\sim 10^{5}\,R_{\rm s}$ ($\sim 1.5\times 10^{4}\,R_{\rm s}$) with DECIGO (LISA).
Furthermore, for tight orbits around low mass black holes we can extract these quantities with sub-percent precision. All results in this analysis assume $\cos\vartheta$ to be 0.9. The angular location of the binary in the orbital plane at coalescence is a nuisance parameter that needs to be marginalized out when inferring other more astrophysically relevant parameters such as the radial location of the binary in the environment, and the environment’s potential profile. Given the spherically symmetric nature of the potentials considered in this work, we do not expect the choice of this value to drastically change the inference of environment and orbit parameters. Nevertheless, to ascertain this, we show corner plots of the recovery of kinematic and environmental parameters in Figure 8 of the Appendix D. We assume a $30-30\,{\rm M}_{\odot}$ binary in circular orbit of radius $25\,R_{\rm s}$ around a $10^{6}\,{\rm M}_{\odot}$ SMBH, at $100\,{\rm Mpc}$, in ET, with $\cos\vartheta=0.5$. As with the $\cos\vartheta=0.9$ case, all of the parameters are recovered well within the posterior.
Figure 4: The relative errors in the semi-major axis and eccentricity of the outer orbit over a grid of SMBH masses $M_{\mathrm{SMBH}}$ and semi-major axis $R$ for the binary systems and detector configurations considered in Figure 3 of the paper. Top Panels: Relative error in the semi-major axis. The ‘solid’ lines are the lines of constant $\frac{\Delta R}{R}$. Bottom Panels: Relative error in the eccentricity. The ‘dashed’ lines are the lines of constant $\frac{\Delta e}{e}$. The pink regions, the blue lines, and the arrows have the same meaning as in Figure 2.
III.2 Elliptical Orbit around an SMBH
For binaries in an elliptical orbit around an SMBH, the kinematic parameters along our LOS depend on $M_{\rm SMBH}$, semi-major axis ($R$), eccentricity ($e$), and $x\equiv R/r$ where $r$ is distance of the CoM of the binary from the gravitating center, at $t_{\rm c}$.
Therefore, together with LOSA, LOSJ, and LOSS, we will need one more derivative of LOSV: LOS crackle (LOSC), to uniquely determine these quantities (see Appendix C.2 for the expressions of kinematic parameters).
@@ III.2 Elliptical Orbit around an SMBH
    With LIGO at A+ sensitivity, we can successfully extract the mass of a $10^{5}\,{\rm M}_{\odot}$ SMBH and localise a $5-1.4\,{\rm M}_{\odot}$ binary having optimal SNR $\sim 145$ corresponding to $D_{\rm L}=100\,{\rm Mpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 250\,R_{\rm s}$.
@@
    With ET, we can successfully extract the mass of the same SMBH and localise a $30-30\,{\rm M}_{\odot}$ binary having optimal SNR $\sim 8248$ corresponding to $D_{\rm L}=100\,{\rm Mpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 500\,R_{\rm s}$, and within a precision of $\sim 20\%$ up to $R\sim 200\,R_{\rm s}$.
@@
    We can successfully infer the masses of SMBHs as massive as $10^{8}\,{\rm M}_{\odot}$, and localise a $100-100\,{\rm M}_{\odot}$ binary having optimal SNRs $\sim 10782$ and $\sim 14$, corresponding to $D_{\rm L}=1\,{\rm Gpc}$, in an elliptical orbit of eccentricity $0.5$, up to a semi-major axis $R\sim 5\times 10^{4}\,R_{\rm s}$ and $\sim 3\times 10^{4}\,R_{\rm s}$ with DECIGO and LISA, respectively.
The chosen value of $x$ is a random draw from a uniform distribution between $\frac{1}{1-e},\frac{1}{1+e}$. We also find that, for certain portions of the parameter space, we can achieve sub-percent precision in DECIGO (see lower left regions of the parameter space in Figure 3).
Figure 4 shows the relative errors in the semi-major axis ($R$) and eccentricity ($e$) of the outer orbit over a grid of SMBH masses $M_{\mathrm{SMBH}}$ and $R$, with $e=0.5$. The binary systems and detector configurations considered are the same as those in Figure 3. We can successfully extract $R$ and $e$, around a $10^{5}\,{\rm M}_{\odot}$ SMBH, out to $R\sim 400\,R_{\rm s}$ with A+, and $R\sim 500\,R_{\rm s}$ with ET. We can do the same around a $10^{6}\,{\rm M}_{\odot}$ SMBH out to $R\sim 7\times 10^{5}\,R_{\rm s}$ with DECIGO, and $R\sim 10^{6}\,R_{\rm s}$ with LISA. Furthermore, it is possible to extract these quantities within a sub-percent level of accuracy in certain portions of the parameter space (the lower left regions of the panels of Figure 4).
Note that we have verified that the CBCs considered in this paper will not be tidally disrupted by the SMBH, as the minimum semimajor axis required to avoid tidal disruption is much smaller than the values considered here.
III.3 Effective Keplerian Orbit in BW potential
For binaries in NSCs with a BW potential profile, the kinematic parameters projected along the LOS will depend on $M_{\rm MBH}$, $r$, steepness of the potential profile ($\alpha$), radial ($v_{r}$) and angular ($\Omega$) speeds of the CBC’s CoM at coalescence, and $\cos\vartheta$. Hence, together with LOSA, LOSJ, LOSS, and LOSC, we will need two more derivatives of LOSV to break all degeneracies among the inferred parameters.
@@ III.3 Effective Keplerian Orbit in BW potential
    With DECIGO, we can extract the mass of a $10^{6}\,{\rm M}_{\odot}$ MBH within a precision of $\sim 10\%$ together with $\alpha$, and localise a $30-30\,{\rm M}_{\odot}$ binary having optimal SNR $\sim 3951$ corresponding to $D_{\rm L}=1\,{\rm Gpc}$, with a sub-percent precision out to $\sim 3\times 10^{5}\,R_{\rm s}$.
@@
    With LISA, we can extract the mass of a $10^{6}\,{\rm M}_{\odot}$ MBH within a precision of $\sim 25\%$ together with $\alpha$ and localise a $100-100\,{\rm M}_{\odot}$ binary having optimal SNR $\sim 14$ corresponding to $D_{\rm L}=1\,{\rm Gpc}$, with a sub-percent precision out to $\sim 3\times 10^{4}\,R_{\rm s}$.
@@
It is clear from Figure 3 that we can extract $M_{\rm MBH}$ and $r$ with sub-percent precision for certain parts of this parameter space, for both DECIGO and LISA. However, we cannot do so for $\alpha$ with LISA, for the fiducial CBC we consider. This can be understood as follows. Larger magnitudes of kinematic parameters increase our ability to constrain environment parameters. Getting closer to the MBH does indeed increase these magnitudes; however, the MBH potential starts to dominate over the potential with exponent $\alpha$. As a result, the precision of $\alpha$ gets suppressed. There is therefore a sweet spot in the $M_{\rm MBH}-r$ parameter space where the precision on the inference of $\alpha$ is optimized. This spot changes, depending on intrinsic parameters of the CBC, including its orbital parameters. Figure 3 does not clearly show this sweet spot because it lies within the highly-relativistic domain where our analysis is not carried out.
As an illustration, in Figure 9 of Appendix D, we have shown the recoveries of the mass of the central massive black hole $M_{\mathrm{MBH}}$, the radial location of the CBC in the outer orbit $r$, and the steepness of the gravitational potential $\alpha$, for a $30-30\,\mathrm{M}_{\odot}$ binary at $1\,\mathrm{Gpc}$, in DECIGO.
@@
Figure 5: Effective Keplerian outer-orbit in an NSC with a BW profile: The precision of the extraction of parameters over a grid of MBH masses $M_{\rm MBH}$ and merger distance $r$ for a binary system of component masses (from left to right): $30-30\,{\rm M}_{\odot}$ and $100-100\,{\rm M}_{\odot}$ at $1{\rm Gpc}$ in DECIGO and LISA, respectively, for $\alpha=1.75$ and $\cos\vartheta=0.9$ with radial and tangential velocities: $v_{r}=250\,{\rm km/s}$ and $r\Omega=100\,{\rm km/s}$ [^3] These numbers are consistent with observed proper motions of stars in the Galactic NSC (see Figure 6 of Schödel et al. (2009))., respectively. Top Panel: Relative error in the mass of the MBH. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{\rm MBH}}{M_{\rm MBH}}$. Middle Panel: Relative error in the distance of the binary from the NSC’s center. The ‘dashed’ lines are the lines of constant $\frac{\Delta r}{r}$. Bottom Panel: Relative error in the steepness parameter $\alpha$. The ‘dotted’ lines are the lines of constant $\frac{\Delta\alpha}{\alpha}$. The pink regions, the blue lines, and the arrows have the same meaning as in Figure 2.
III.4 Effective Keplerian Orbit in Plummer potential
@@
For binaries in GCs with a Plummer potential Plummer (1911); Aarseth et al. (1974) profile, kinematic parameters along our LOS will depend on the total mass of the cluster ($M_{0}$), $r$, Plummer radius ($a_{\mathrm{p}}$), $v_{r}$, $\Omega$, and $\cos\vartheta$. Therefore, similar to the BW case, we will need six derivatives of LOSV to uniquely determine parameters. However, the Plummer profile being shallower than the BW profile, recovering LOSC and higher time derivatives of LOSV is not possible for the range of parameters we consider. Hence, we only evaluate the LOSA, LOSJ, and LOSS (see C.4 for the expressions) when constructing the GW phase modulations.
@@
Figure 6 shows the relative errors in the measurements of $M_{0}$ (top panel), $r$ (middle panel), and $a_{\mathrm{p}}$ (bottom panel) over a grid of $M_{0}$ and $r$ for $a_{\rm p}=0.1\,{\rm pc}$. We find that with DECIGO, we can successfully extract $M_{0}$, $a_{\rm p}$, and localise a $30-30\,\rm M_{\odot}$ binary having optimal SNR $\sim 39513$ corresponding to $D_{\rm L}=100\,{\rm Mpc}$, for total cluster masses greater than $\sim 3\times 10^{5}\,{\rm M}_{\odot}$.
Even though we can successfully constrain $M_{0}$, $r$, and $a_{\mathrm{p}}$ out to $r\sim 0.4\mathrm{pc}$, it still requires the cluster to be as heavy as $10^{7}\,{\rm M}_{\odot}$ which is well above the most massive known GC D’Souza and Rix (2013). With LISA, LOSJ is only measurable for a very limited portion of the parameter space we consider. Therefore, we do not show the corresponding results.
@@
Figure 6: Effective Keplerian outer-orbit in a GC with Plummer profile: The precision of the extraction of parameters over a grid of cluster masses $M_{0}$ and merger distance $r$ for a binary system of component masses $30-30\,{\rm M}_{\odot}$ at $100\mathrm{Mpc}$ in DECIGO with $\cos\vartheta=0.25$. Left Panel: Relative error in the total mass of the cluster. The ‘solid’ lines are the lines of constant $\frac{\Delta M_{0}}{M_{0}}$. Middle Panel: Relative error in the distance of the binary from the GC’s center. The ‘dashed’ lines are the lines of constant $\frac{\Delta r}{r}$. Right Panel: Relative error in the Plummer radius where $a_{\rm p}=0.1\,{\rm pc}$. The ‘dotted’ lines are the lines of constant $\frac{\Delta a_{\mathrm{p}}}{a_{\mathrm{p}}}$. The pink regions and the arrows have the same meaning as in Figure 2.
III.5 Stealth Bias
An inherent assumption of our analysis is that the correct model to be fit to the data is known. This will not be true in general. Moreover, use of the wrong model could lead to stealth biases. We show this in Figure 11 of the Appendix D where a $100-100\,{\rm M_{\odot}}$ binary at $1\,{\rm Gpc}$ in an eccentric outer orbit of semi-major axis $R=5\times 10^{4}\,R_{\rm s}$ around a $10^{6}\,{\rm M_{\odot}}$ SMBH, detected in DECIGO, is considered. For eccentricities (of the outer orbits) as low as $0.01$, fitting a circular orbit will result in a stealth bias of one of the outer orbit parameters. Nevertheless, it is possible to use Bayesian model selection first to determine which orbit is preferred: circular or eccentric. For this, we set $\cos\vartheta=-e$ in Eq. 43 to set $x=1$ (i.e. $r=R$) so that we have only one extra parameter eccentricity $e$ in the true model. This allows us to use the Savage-Dickey density ratio to evaluate a Bayes Factor (see Zimmerman et al. (2019) and the references therein). Following Eq. (5) of Zimmerman et al. (2019), we define the Bayes factor in favor of a non-zero eccentricity as:
[…]
where $\mu_{e}$ is the median of the posterior of eccentricity, $\sigma$ is the standard deviation of the same, while $\Delta\lambda$ is the prior on the eccentricity which we take to be uniform between 0 and 1. As shown in Figure 7, we find strong preferences for eccentric orbits, even for eccentricities as low as $e=0.01$, with Bayes factors increasing drastically for larger eccentricities.
[…]
Figure 7: The variation of ($\log_{10}$) Bayes Factor $\mathcal{B}$ in favor of a non-zero eccentricity with the eccentricity of the outer orbit $e$ for the scenario considered in Figure 11. The inset plot shows the variation of the same at low eccentricities of the outer orbit. Even for $e=0.01$, $\mathcal{B}\sim 100$.
IV Discussion